OpenStudy (anonymous):

What is the solution set for |2x – 4| = 10?

OpenStudy (anonymous):

We need two equations. First one is 2x - 4 = -10 What is the other equation?

OpenStudy (ksaimouli):

7

OpenStudy (saifoo.khan):

|2x – 4| = 10 2x – 4 = 10 and 2x – 4 = - 10 Solve both! :)

OpenStudy (ksaimouli):

and-7

OpenStudy (anonymous):